Suppose a random variable is normally distributed. The probabilities for a85 and 142 are 10% and 65%, respectively. Find the mean and standard deviation to the nearest whole number.

1
Expert's answer
2021-12-07T09:49:11-0500
"P(X<85)=P(Z<\dfrac{85-\mu}{\sigma})=0.1"


"\dfrac{85-\mu}{\sigma}\approx-1.2815516"


"P(X<142)=P(Z<\dfrac{142-\mu}{\sigma})=0.65"


"\dfrac{142-\mu}{\sigma}\approx0.385320"

"\dfrac{85-\mu}{-1.2815516}=\dfrac{142-\mu}{0.385320}"

"\mu=\dfrac{85(0.385320)+142(1.2815516)}{0.385320+1.2815516}"

"\mu\approx128.8237"

"\sigma=\dfrac{142-128.8237}{0.385320}"

"\sigma=34.1957"

Answer:

"\mu=129, \sigma=34"